OpenStudy (anonymous):

Jason ran for 1.5 hours with a 2 mph wind at his back. It took him 2.5 hours to get home with the same wind in his face. How would Jason have run if not for the wind? a. 2mph b. 6 mph c. 8 mph d. 4 mph

OpenStudy (anonymous):

d=vt He ran the same distance back and forth. (v+2)(1.5)=(v-2)(2.5) 1.5v+3=2.5v-5 v=8

OpenStudy (cwrw238):

let his speed without any wind be x mph x + 2 = d /1.5 where d = distance run x - 2 = d /2.5 1.5x + 3 = d 2.5x - 5 = d subtract x - 8 = 0 so x = 8 8 mph